# EpiCache: Episodic KV Cache Management for Long Conversational Question Answering

### v1

@@ 1 Introduction
Large language models (LLMs) (Brown et al., 2020; Yang et al., 2025; Touvron et al., 2023; Jiang et al., 2023) have significantly extended their context lengths, with LLM-based chat assistants now capable of processing hundreds of thousands to millions of tokens (Reid et al., 2024; Meta, 2025). This capability enables assistants to leverage extensive dialogue histories when generating responses, producing personalized and contextually coherent outputs (OpenAI, 2024; Anthropic, 2024), which are central requirements for real-world conversational AI applications.
@@
Prior work has attempted to mitigate the growing memory footprint of KV caches in long conversations through various compression techniques (Zhang et al., 2023; Li et al., 2024; Cai et al., 2025). Yet two major limitations remain under memory-constrained settings. First, most methods apply compression after full context prefill (post-prefill), causing peak memory usage to scale linearly with input length. Second, query-dependent eviction (Li et al., 2024) optimizes for the current question but ties the cache closely to it, neglecting information needed for future queries and degrading accuracy in multi-turn conversations.
We propose EpiCache, a training-free KV cache management framework that enforces a constant memory footprint through block-wise prefill. After each block, less critical KV entries are evicted, ensuring memory consumption remains bounded. Beyond block prefill, EpiCache incorporates an episodic clustering method inspired by conversation segmentation studies (Joty et al., 2013; Galley et al., 2003). Specifically, we apply semantic clustering to group conversation history into coherent episodes, and perform episodic KV cache compression, yielding topic-specific caches within the fixed budget.
@@ 2 Background
We begin by formalizing Long Conversational Qustion-Answering (LongConvQA) in Section 2.1. We then discuss memory-constrained KV cache management in Section 2.2, comparing post- and block-prefill eviction and discussing the resulting accuracy-memory trade-offs. Finally, in Section 2.3, we review attention-guided cache compression with patched prompts and present analyses that motivate our method, EpiCache.
@@ 2.1 Long Conversational QA Formulation
This formulation serves as an evaluation of multi-turn conversational accuracy, where multiple query-answer pairs are grounded in the same dialogue history.
@@ 2.2 KV Cache Management: Post Prefill vs Block Prefill
To bound memory growth, Block Prefill Eviction (Kim et al., 2024; Corallo & Papotti, 2024; Park et al., 2025) processes the input in a block-wise way, handling one segment at a time under a fixed budget $M$. Each step adds $M_{\text{block}}$ tokens, after which eviction reduces KV cache entries back to $M$. For example, in Figure 1(b), the budget is $M=1$, and each block adds $M_{\text{block}}=3$ tokens that are evicted to $M$. This design ensures the cache never exceeds $M+M_{\text{block}}$, keeping peak GPU memory usage flat with input length as highlighted in Figure 1(c) top.
@@ 2.3 Attention-guided KV Cache Compression
To address the accuracy degradation of block prefill eviction, prior work employs attention-based token scoring with patched prompt. Here, token importance is quantified by the cross-attention it receives from query tokens: $\mathrm{Attn}(x_{t}\!\rightarrow\!x_{i})$ which denotes the attention weight from a query token $x_{t}$ to a key token $x_{i}$. Tokens that receive higher attention from queries are considered more important, while those with lower scores are evicted to satisfy the memory budget $M$.
To provide guidance for cache eviction, the patched prompt strategy (Kim et al., 2024; Bhaskar et al., 2025) appends an auxiliary prompt of length $p$ after each block ending at token $n$. These queries $x_{n+1},\dots,x_{n+p}$} attend back to the preceding block tokens $x_{1},\dots,x_{n}$}, as illustrated in Figure 1(b). The resulting importance score $s_{i}$ of token $i$ is aggregated either by averaging or by taking the maximum, as defined in Equation 3 (the patched prompt is only used for scoring and not retained in the KV cache; all experiments in this paper adopt the maximum aggregation).
@@
In block prefill, accuracy is highly dependent on the content of the patched prompt. To study this, Figure 2 presents a controlled experiment where we assume oracle access to the future user query, with inserting it as the patched prompt yielding the highest accuracy (Exact-Question) [^1] This strategy is infeasible in LongConvQA since (i) the future queries are unknown at compression time, and (ii) each new query requires a new prefill and compression step (Kim et al., 2025).. Since the dialogue history $\mathcal{H}$ in Equation 1 consists of question-answer turns, it offers an opportunity to approximate the future query with semantically related utterances. To test this idea, we embed both user queries $q_{1},\dots,q_{N_{q}}$ and conversation utterances $u_{1},\dots,u_{N_{u}}$ into a shared embedding space, compute semantic similarity scores, and construct patched prompts using the top-ranked segments.
@@
As expected, the results confirm that conversation utterances semantically aligned with the future question can serve as effective proxies, with the closest 10% achieving accuracy nearly matching the Exact-Question case in Figure 2. Accuracy then declines as the selected segments become less related, showing that the degree of semantic alignment directly determines the quality of answers generated from the compressed KV cache. These findings narrow our objective: during block prefill eviction, the central challenge is to identify a patched prompt that best approximates unseen future questions. To this end, we employ unsupervised methods to discover dialogue segments aligned with future queries, motivating the clustering-based approach that will be introduced in Section 3.1.
@@ 3.1 Episodic KV Cache Management with Conversation Clustering
EpiCache consists of three stages as illustrated in Figure 3. Stage 1 (Clustering) is performed offline, where the conversation history is clustered into topical episodes, and a representative segment is identified for each episode. Stage 2 (Prefilling) leverages representative segments as a patched prompt to guide block prefill eviction, building episode-specific KV cache while keeping GPU memory bounded. Stage 3 (Decoding) embeds the incoming query, matches it to the most relevant episode, and retrieves the corresponding episodic KV cache for answer generation.
@@
For clustering conversation, we first divide the dialogue history $\mathcal{H}$ into segments of $w_{\text{embed}}$ utterances, denoted as $\mathcal{H}={S_{1},\dots,S_{K}}$.
@@
Each segment $S_{k}$ is encoded with a sentence encoders (Reimers & Gurevych, 2019) $f_{\text{embed}}$ into a semantic vector representation $\mathbf{e}_{k}\in\mathbb{R}^{d}$, which captures the conversational content of the segment. We then apply K-Means clustering $\mathcal{C}(\cdot)$ to the embeddings $\{\mathbf{e}_{k}\}_{k=1}^{K}$, as illustrated in Figure 3(a):
@@
(c) Episodic KVs based Decoding
@@
We then identify the medoid segment—i.e., the conversation segment in each clusters whose embedding is most similar to the centroid in terms of semantic similarity. The medoid segment, as the representative of the cluster, contains multiple turns from both speakers and is used as the patched prompt in the subsequent block prefill eviction step.
@@
As discussed in Section 2.3, patched prompts guide cache eviction toward retaining tokens relevant to the prompt content. Building on this insight, EpiCache uses the medoid segment of each episode as the patched prompt, enabling it to collect episode-specific KV entries under the memory budget. For each episode $e\in\{1,\dots,E\}$, we perform block prefill eviction over entire context, appending its patched prompt $\mathcal{P}_{e}$ after each block (Figure 3(b)). Attention scores are then computed with $\mathcal{P}_{e}$, and the top $M$ tokens are retained to form an episode-specific cache $C_{\mathrm{KV}}^{(e)}$. Finally, all episodic caches are collected into $\mathbb{B}=\{C_{\mathrm{KV}}^{(1)},\dots,C_{\mathrm{KV}}^{(E)}\}$ and stored offline for later retrieval.
@@
As illustrated in Figure 3(c), the framework retrieves the corresponding episodic cache $C_{\mathrm{KV}}^{(e^{\dagger})}$ from $\mathbb{B}$ and conditions generation on it: $f_{\text{LM}}(q_{i}\mid C_{\mathrm{KV}}^{(e^{\dagger})},)$. This design enables query-specific retrieval from compressed episodic caches while keeping cache size bounded under memory budget $M$. This query-to-episode matching process incurs overhead from embedding, centroid matching, and cache retrieval, which will be analyzed in Section 4.4.
@@
Figure 4: Layer-wise Sensitivity Analysis and KV Budget Allocation. (a) Key state cosine similarity across normalized layer positions. (b) KL divergence is measured between block prefill (M=4M{=}4K) and full KV answer predictions, with uniform allocation as the baseline. Per-sample divergence shifts are shown when applying three allocation strategies—sensitivity-aware, PyramidKV, and retrieval head profiling-on the Realtalk benchmark with Qwen2.5-7B. (c) LongMemEval accuracy comparison across allocation methods using Qwen2.5-7B.
@@ 3.2 Sensitivity-aware Layer-wise KV Budget Allocation
To quantify the deviation caused by block prefill eviction, we introduce a custom masking scheme. Each transformer layer is represented as a simplified function $f$ that takes the previous layer’s output $X^{(\ell-1)}\in\mathbb{R}^{N\times d}$, where $N$ is the sequence length and $d$ the hidden dimension, and produces the $l_{\text{th}}$ layer output $X^{\ell}$ under a given mask $\mathcal{M}$: $X^{\ell}=f\left(X^{\ell-1},\mathcal{M}\right).$
While $\mathcal{M}$ is the causal mask, we replace it with a custom mask $\mathcal{M}^{\prime}$ that uses a budget $M$, while attending sink tokens and the most recent tokens (Xiao et al., 2024). This simulates block prefill eviction in a single forward pass, enabling direct measurement of its effect on layer representations.
@@
We quantify per-layer impact by comparing Key states produced under the causal mask $\mathcal{M}$ and the custom mask $\mathcal{M}^{\prime}$. [^2] Further details regarding the rationale for using Key states deviation are provided in Section C.1. For each layer $\ell$, the forward pass under each mask produces:
@@
where $K^{\ell}_{\text{full}}$ and $K^{\ell}_{\text{block}}$ are the $l$-th layer Key states computed under $\mathcal{M}$ and $\mathcal{M}^{\prime}$, respectively. We then define layer sensitivity as the average cosine similarity between the two sets of Key vectors across heads and tokens:
$\sigma_{\ell}=\tfrac{1}{HN}\sum_{h=1}^{H}\sum_{i=1}^{N}\cos\!\big(k^{(\ell,h)}_{\text{full},i},\,k^{(\ell,h)}_{\text{block},i}\big),\quad s_{\ell}=1-\sigma_{\ell}$ (9)
Empirically, $\sigma_{\ell}$ exhibits large variation across layers yet remains consistent across different inputs in Figure 4(a) (the shadowed regions denote input variance), indicating that sensitivity is model-dependent rather than input-dependent. We define $s_{\ell}$ as the sensitivity score for layer $\ell$.
@@
Figure 5: LongConvQA Evaluation Results (Realtalk, LoCoMo, and LongMemEval) results with fixed KV cache budget size-$M$ across four LLMs. The number of episodes (clusters) fixed to E=4E{=}4 in all experiments. The average full KV lengths of the three benchmarks are 26K, 21K, and 21K.
@@ 4.1 Setup
EpiCache clusters the conversation history $\mathcal{H}$ into episodic segments and applies patched-prompt KV cache compression, with per-layer budgets allocated by sensitivity measurements. We use E=4E{=}4 episodes and Qwen3-0.6B (Zhang et al., 2025) as the embedding model. Layer sensitivities are profiled once on a BookSum (Kryściński et al., 2022) sample for each LLM and reused across all experiments, with the sharpness hyper-parameter set to $\alpha{=}1.1$ for the LLaMA series and $\alpha{=}1.3$ for the Qwen series. The overall process is detailed in Algorithm 1, while further setup specifications are provided in Section A.3.
@@ 4.2 Main Evaluation Results
Figure 5 shows results on the three LongConvQA benchmarks—Realtalk, LoCoMo, and LongMemEval—under block prefill eviction with varying cache budget $M$. StreamingLLM consistently shows the lowest accuracy, while patched-prompt methods (SnapKV, InfiniPot, KVzip) and Key states similarity-based KeyDiff suffer significant accuracy degradation compared to full KV. EpiCache achieves significantly higher scores across all models and benchmakrs, exceeding baselines by around 20 scores under tightly compressed budgets (M=M{=}2-4K), and approaching full KV performance at budgets (M=M{=}6-8K). These consistent improvements across the three benchmarks highlight the effectiveness of episodic KV cache compression, along with sensitivity-aware budget allocation, in preserving conversational context under fixed memory constraints. Detailed results including sub-task accuracy are provided in Appendix D.
Figure 6: Memory Scalability up to 100K Context. Conversation histories between user and LLMs scaled to 100K tokens across four LLMs with LongMemEval. Comparison of InfiniPot and KVzip (M=6M{=}6K) with EpiCache (4 episodes, M=6M{=}6K–24K).
@@
Figure 6 evaluates LongConvQA under extended conversation lengths, scaling up to 100K tokens. [^3] LongMemEval (Wu et al., 2025a) supports stacking conversation sessions with associated QA pairs, allowing conversation histories to be constructed at custom lengths. Open-source LLMs exhibit declining QA performance as context length grows to 100K, as reported in Wu et al. (2025a), and the performance gap between full KV and cache compression methods (KVzip, InfiniPot) becomes increasingly pronounced. EpiCache delivers higher accuracy than baselines at the same memory budget across all context lengths, and as the KV cache budget increases, its accuracy steadily approaches full KV, demonstrating the memory scalability of our approach.

### v2

@@ 1 Introduction
Large language models (LLMs) (Brown et al., 2020; Yang et al., 2025; Touvron et al., 2023; Jiang et al., 2023) have significantly extended their context lengths, with LLM-based AI assistants now capable of processing millions of tokens (Reid et al., 2024; Meta, 2025). This capability enables assistants to leverage extensive dialogue histories when generating responses, producing personalized and contextually coherent outputs (OpenAI, 2024; Anthropic, 2024), which are central requirements for conversational AI applications (Fu et al., 2022).
@@
Prior work has attempted to mitigate the growing memory footprint of KV caches in long conversations through various compression techniques (Zhang et al., 2023; Li et al., 2024; Cai et al., 2025). Yet two major limitations remain under memory-constrained settings. First, most methods apply compression after prefilling the entire input context (post-prefill), causing peak memory usage that scales linearly with input length. Second, query-dependent eviction (Li et al., 2024) retains cache entries by focusing on the current query, and neglects information needed for future queries, thereby degrading accuracy in multi-turn conversations.
We propose EpiCache, a training-free KV cache management framework that enforces a constant memory footprint through block-wise prefill. After processing each block, we evict less critical KV entries to free space for the next block, ensuring memory consumption remains bounded. Based on block prefill, EpiCache incorporates an episodic clustering method inspired by conversation segmentation studies (Joty et al., 2013; Galley et al., 2003). Specifically, we apply semantic clustering to group conversation history into coherent episodes, and perform episodic KV cache compression, yielding topic-specific caches while using constrained memory.
@@ 2 Background
We begin by formalizing Long Conversational Question-Answering (LongConvQA) in Section 2.1. We then discuss memory-constrained KV cache management in Section 2.2, comparing post- and block-prefill eviction and discussing the resulting accuracy-memory trade-offs. Finally, in Section 2.3, we review attention-guided cache compression with patched prompts and present analyses that motivate our method, EpiCache.
@@ 2.1 Long Conversational QA Formulation
This formulation serves as an evaluation of multi-turn conversation accuracy, where multiple query-answer pairs are grounded in the same dialogue history.
@@ 2.2 KV Cache Management: Post Prefill vs Block Prefill
To bound memory growth, Block Prefill Eviction (Kim et al., 2024; Corallo & Papotti, 2024; Park et al., 2025) processes the input in a block-wise way, handling one segment at a time under a fixed budget $M$. Each step adds $M_{\text{block}}$ tokens, after which eviction reduces KV cache entries back to $M$. For example, in Figure 1(b), the budget is $M=1$, and each block adds $M_{\text{block}}=3$ tokens that are evicted to $M$. This design ensures the number of cache entries never exceeds $M+M_{\text{block}}$, keeping peak GPU memory usage flat with input length as highlighted in Figure 1(c) top.
@@ 2.3 Attention-guided KV Cache Compression
To address the accuracy degradation of block prefill eviction, prior work employs attention-based token scoring with a patched prompt. Here, token importance is quantified by the cross-attention it receives from query tokens: $\mathrm{Attn}(x_{t}\!\rightarrow\!x_{i})$ which denotes the attention weight from a query token $x_{t}$ to a key token $x_{i}$. Tokens that receive higher attention from queries are considered more important, while those with lower scores are evicted to satisfy the memory budget $M$.
To provide guidance for cache eviction, the patched prompt strategy (Kim et al., 2024) appends an auxiliary prompt of length $p$ after each block ending at token $n$. These queries $x_{n+1},\dots,x_{n+p}$} attend back to the preceding block tokens $x_{1},\dots,x_{n}$}, as illustrated in Figure 1(b). The resulting importance score $s_{i}$ of token $i$ is aggregated either by averaging or by taking the maximum, as defined in Equation 3. The patched prompt is used for scoring and not retained in the KV cache, and all experiments in this paper adopt the maximum aggregation.
@@
In block prefill, accuracy is highly dependent on the content of the patched prompt. To study this, Figure 2 presents a controlled experiment where we assume oracle access to the future user query, with inserting it as the patched prompt yielding the highest accuracy (Exact-Question) [^1] This strategy is infeasible in LongConvQA since (i) the future queries are unknown at compression time, and (ii) each new query requires re-prefill of entire context (Kim et al., 2025).. Since the dialogue history $\mathcal{H}$ in Equation 1 consists of question-answer turns, it offers an opportunity to approximate the future query with semantically related turns. To test this idea, we embed both user queries $q_{1},\dots,q_{N_{q}}$ and conversation utterances $u_{1},\dots,u_{N_{u}}$ into a shared embedding space, compute semantic similarity scores, and construct patched prompts using the top-ranked multi-turns.
@@
As expected, the results confirm that conversation utterances semantically aligned with the future question can serve as effective proxies, with the closest 10% achieving accuracy nearly matching the Exact-Question case in Figure 2. Accuracy then declines as the selected segments become less related, showing that the degree of semantic alignment directly determines the quality of answers generated from the compressed KV cache. These findings narrow our objective: during block prefill eviction, the central challenge is to identify a patched prompt that best approximates unseen questions without any supervision from future conversation. To this end, we employ unsupervised clustering methods to discover dialogue segments aligned with future queries, which will be introduced in Section 3.1.
@@ 3.1 Episodic KV Cache Management with Conversation Clustering
Conversations can be segmented into coherent episodes (Galley et al., 2003; Sieber & Krenn, 2010), and subsequent utterances are naturally grounded in prior episodes. This episodic nature drives the core insight of EpiCache: by clustering dialogue into episodes and constructing episode-specific caches, we can match an incoming query to the most relevant cache for accurate answer generation. As illustrated in Figure 3, this process unfolds in three stages—lightweight conversation clustering, episodic KV cache construction, and decoding with query matching.
@@
For clustering conversation, we first divide the raw dialogue history $\mathcal{H}$ into segments of $w_{\text{embed}}$ utterances, denoted as $\mathcal{H}={S_{1},\dots,S_{K}}$.
@@
Each segment $S_{k}$ is encoded with a sentence encoders (Reimers & Gurevych, 2019) $f_{\text{embed}}$ into a vector embeddings $\mathbf{e}_{k}\in\mathbb{R}^{d}$, capturing the segment’s semantics. We then apply K-Means clustering $\mathcal{C}(\cdot)$ to the embeddings $\{\mathbf{e}_{k}\}_{k=1}^{K}$, as illustrated in Figure 3(a):
@@
(c) Decoding with Episodic KVs
@@
We then identify the medoid segment—i.e., the conversation segment in each cluster whose embedding is closest to the centroid in terms of semantic similarity. The medoid segment, as the representative of the cluster, contains multiple turns from both speakers and is used as the patched prompt in the subsequent block prefill eviction step.
@@
As discussed in Section 2.3, patched prompts guide cache eviction toward retaining tokens relevant to the prompt content. Building on this insight, EpiCache uses the medoid segment of each episode as the patched prompt, thereby collecting episode-specific KV entries under the memory budget.
For each episode $e\in\{1,\dots,E\}$, we perform block prefill eviction over entire context, appending its patched prompt $\mathcal{P}_{e}$ after each block (Figure 3(b)). Attention scores are then computed with $\mathcal{P}_{e}$, and the top $M$ tokens are retained to form an episode-specific cache $C_{\mathrm{KV}}^{(e)}$. Finally, all episodic caches are collected into $\mathbb{B}=\{C_{\mathrm{KV}}^{(1)},\dots,C_{\mathrm{KV}}^{(E)}\}$ and stored offline for later retrieval.
@@
As illustrated in Figure 3(c), the framework retrieves the corresponding episodic cache $C_{\mathrm{KV}}^{(e^{\dagger})}$ from $\mathbb{B}$ and conditions generation on it: $f_{\text{LM}}(q_{i}\mid C_{\mathrm{KV}}^{(e^{\dagger})},)$. This design enables query-specific retrieval from episodic caches while keeping cache size bounded under memory budget $M$. This query-to-episode matching introduces overhead from embedding, matching, and retrieval; however, the cost is minor compared to decoding and is further examined in Section 4.4.
@@
Figure 4: Layer-wise Sensitivity Analysis and KV Budget Allocation. (a) Key states cosine similarity across normalized layer positions. (b) KL divergence is measured between block prefill (M=4M{=}4K) and full KV answer predictions, with uniform allocation as the baseline. Per-sample KL divergence shifts are shown when applying three allocation strategies—sensitivity-aware, PyramidKV, and retrieval head profiling-on the Realtalk benchmark with Qwen2.5-7B. (c) Realtalk accuracy comparison across budget allocation methods using Qwen2.5-7B.
@@ 3.2 Sensitivity-aware Layer-wise KV Budget Allocation
To quantify the deviation caused by block prefill eviction, we introduce a custom masking scheme. Each transformer layer is represented as a function $f$ that takes the previous layer’s output $X^{(\ell-1)}\in\mathbb{R}^{N\times d}$. Here, $N$ denotes the sequence length and $d$ the hidden dimension. The function produces the $\ell_{\text{th}}$ layer output $X^{\ell}=f\left(X^{\ell-1},\mathcal{M}\right)$, where $\mathcal{M}$ is the standard causal mask.
We replace $\mathcal{M}$ with a custom mask $\mathcal{M}^{\prime}$ that enforces a budget $M$, attending to sink tokens and the most recent tokens (Xiao et al., 2024). This design follows static compression methods, allowing us to simulate block prefill eviction in a single forward pass and directly measure its effect on layer representations.
@@
We quantify the per-layer impact of block prefill eviction by comparing Key states [^2] Further details regarding the rationale for using Key states deviation are provided in Section C.1. produced under the causal mask $\mathcal{M}$ and the custom mask $\mathcal{M}^{\prime}$. For each layer $\ell$, the forward pass under each mask produces:
@@
where $K^{\ell}_{\text{full}}$ and $K^{\ell}_{\text{block}}$ are the $l$-th layer Key states computed under $\mathcal{M}$ and $\mathcal{M}^{\prime}$, respectively. We then define layer sensitivity as the average cosine similarity between the two sets of Key vectors across attention heads and input tokens:
$\sigma_{\ell}=\tfrac{1}{HN}\sum_{h=1}^{H}\sum_{i=1}^{N}\cos\!\big(k^{(\ell,h)}_{\text{full},i},\,k^{(\ell,h)}_{\text{block},i}\big)$ (9)
Empirically, $\sigma_{\ell}$ exhibits large variation across layers yet remains consistent across different inputs in Figure 4(a) (shadowed regions denote input variance), indicating that sensitivity is model-dependent rather than input-dependent. We define $s_{\ell}=1-\sigma_{\ell}$ as the sensitivity score for layer $\ell$.
@@
Figure 5: LongConvQA Evaluation Results (Realtalk, LoCoMo, and LongMemEval) results with fixed KV cache budget size-$M$ across four LLMs. The number of episodes (clusters) fixed to E=4E{=}4 in all experiments. The average full KV lengths of the three benchmarks are 26K, 21K, and 20K.
@@ 4.1 Setup
EpiCache clusters the conversation history $\mathcal{H}$ into episodic segments and performs cache compression with patched prompts, while allocating per-layer budgets based on sensitivity measurements. We use E=4E{=}4 episodes and Qwen3-0.6B (Zhang et al., 2025) as the embedding model. Layer sensitivities are profiled once on a BookSum (Kryściński et al., 2022) sample for each LLM and reused across all experiments, with the sharpness hyper-parameter set to $\alpha{=}1.1$ for the LLaMA series and $\alpha{=}1.3$ for the Qwen series. The overall process is detailed in Algorithm 1, while further setup specifications are provided in Section A.3.
@@ 4.2 Main Evaluation Results
Figure 5 shows results on the three LongConvQA benchmarks—Realtalk, LoCoMo, and LongMemEval—under block prefill eviction with varying cache budget $M$. StreamingLLM consistently shows the lowest accuracy, while patched-prompt methods (SnapKV, InfiniPot, KVzip) and Key states similarity-based KeyDiff suffer significant accuracy degradation compared to full KV. EpiCache achieves consistently higher scores across all models and benchmakrs, exceeding baselines by up to 20 points under tightly compressed budgets (M=M{=}2–4K), while approaching full KV performance at budgets (M=M{=}6–8K). These consistent improvements across the three benchmarks highlight the effectiveness of episodic KV cache compression, along with sensitivity-aware budget allocation, in preserving dialogue history under fixed memory constraints. Detailed results including sub-task accuracy are provided in Appendix D.
Figure 6: Memory Scalability up to 100K Context. Conversation histories between user and LLM-based assistant scaled to 100K tokens across four LLMs with LongMemEval. Comparison of InfiniPot and KVzip (M=6M{=}6K) with EpiCache (4 episodes, M=6M{=}6K–24K).
@@
Figure 6 evaluates LongConvQA under extended conversation lengths, scaling up to 100K tokens. [^3] LongMemEval (Wu et al., 2025a) supports stacking conversation sessions with associated QA pairs, allowing conversation histories to be constructed at custom lengths. Open-source LLMs exhibit declining QA performance as context length grows to 100K, as reported in Wu et al. (2025a), and the performance gap between full KV and baseline methods (KVzip, InfiniPot) becomes increasingly pronounced. EpiCache delivers higher accuracy than baselines at the same memory budget across all context lengths, and as the KV cache budget increases, its accuracy steadily approaches full KV, demonstrating the memory scalability of our approach.